# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v3

@@ 3.1 Neighborhood bias
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2},$
and, with probability at least $1-c_{1}n^{-c_{2}}$ (for the same type of constants as ZLZ), uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
Moreover, identifying ZLZ’s terms by
@@
one has the ZLZ-style bounds (with (n1)(n-1) in place of $n$):
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
@@
First, (ZLZ algebra; defines $J_{1},J_{2}$ and matches $U,B$). Exactly as in ZLZ, we begin from the decomposition of the squared error into two contributions:
$\Big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\Big)^{2}=\Bigg[\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\Big\{(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\Big\}\Bigg]^{2},$
and the inequality
$(\cdot)^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
is the same inequality as ZLZ equation (1), with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. (See equation (1), supplementary material of Zhang et al. (2017) .)
Second, two-hop concentration under deletion preserves ZLZ inputs. ZLZ’s neighborhood construction and the control of $J_{2}$ rely on uniform control of two-hop quantities of the form $(A^{2}/n)_{ab}-(P^{2}/n)_{ab}$, established in ZLZ equations, (4)-(6) in their supplementary material. By Lemma 3.2, the same bound holds when replacing $A$ by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, every subsequent ZLZ step that uses (4)-(6) remains valid for the graph with the deleted vertex.
Third, Bounding the $J_{1}$ row-average ZLZ does not bound each $J_{1}(i,\ell)$ separately; instead it controls the row-average $(n^{-1}\sum_{\ell}J_{1}(i,\ell))$ via the decomposition in equation (11), together with the bounds (12), (14), and the conclusion (15) in the ZLZ supplementary material. The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, using equation 3.2 to justify the same concentration inputs. Hence
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},$
by the ZLZ chain (11)-(15).
Fourth, ZLZ bounds $(n^{-1}\sum_{\ell}J_{2}(i,\ell))$ using Cauchy-Schwarz and the “claim 2 of their Lemma 2” input, concluding in equation (16). The same reasoning applies here (again by Lemma 3.2 to justify the same concentration inputs), yielding the same order bound:
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
This is exactly ZLZ equation (16) with the leave-one-out objects.
Finally, we have the same claimed row-wise risk bound, matching ZLZ’s stated goal (their inequality right after introducing $\widetilde{P}$, i.e., the display in the supplementary around the start of the proof of Theorem 1). ∎
@@
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Then there exist constants $c_{0},c_{1}>0$ (depending only on the constants in Assumption 3.5 and on $\varepsilon$) such that, for all sufficiently large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2}\quad\Longrightarrow\quad\begin{cases}U_{i},U_{k}\text{ lie in the same block }I_{m},\\ |U_{i}-U_{k}|\leq c_{1}(\log n/n)^{1/4}.\end{cases}$
@@
Write $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$.
We can easily see that if $U_{i}$ and $U_{k}$ lie in different blocks, then by Assumption 3.5(ii),
$\int_{0}^{1}(f_{i}-f_{k})^{2}\;\geq\;\eta^{2}\delta.$
For $c_{0}>0$ small enough and $n$ large, $c_{0}(\log n/n)^{1/2}<\eta^{2}\delta/2$, so the assumed upper bound on $\frac{1}{n}\sum_{\ell}(P_{i\ell}-P_{k\ell})^{2}$ (which concentrates around $\int_{0}^{1}(f_{i}-f_{k})^{2}$ conditional on $(U_{\ell})$) is incompatible with this. Hence $U_{i},U_{k}$ must lie in the same block $I_{m}$.
On the other hand, since $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$, we have $U_{i},U_{k}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ for some $m$, and Assumption 3.5(iii) gives
$\int_{0}^{1}(f_{i}-f_{k})^{2}=\int_{0}^{1}(W(U_{i},x)-W(U_{k},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}.$
Combining this with the assumed upper bound yields
$\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2},$
so
$|U_{i}-U_{k}|\;\leq\;\sqrt{\frac{c_{0}}{\kappa(\varepsilon)}}\,\Big(\frac{\log n}{n}\Big)^{1/4}.$
Set $c_{1}:=\sqrt{c_{0}/\kappa(\varepsilon)}$. ∎
@@
Assume Assumption 3.5. Let $h_{n}\asymp\sqrt{n\log n}$ and define $r_{n}=(\log n/n)^{1/4}$. For $i\neq j$, set
$B^{(-j)}_{ij}=\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}),\qquad P_{ab}=W(U_{a},U_{b}).$
Suppose there exist constants $C_{2},c>0$ such that, with probability at least $1-n^{-c}$, the LOO neighborhoods satisfy
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}\qquad\text{for all }i\neq j\text{ and all }k\in\mathcal{N}_{i}^{(-j)}.$ (3.4)
Then for all $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\;\leq\;C\,r_{n},$
for some constant $C>0$ depending only on $(L,\eta,\delta,\kappa(\varepsilon),C_{2})$.
@@
Fix $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$ and work on the event where (3.4) holds. Define
$\mathcal{B}_{i}=\Big\{k\in\mathcal{N}_{i}^{(-j)}:\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}>M\Big(\frac{\log n}{n}\Big)^{1/2}\Big\},$
for some constant $M>C_{2}$.
Averaging $(\star)$ over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}.$
If $|\mathcal{B}_{i}|\geq 1$, then the left-hand side is at least
$\frac{|\mathcal{B}_{i}|}{|\mathcal{N}_{i}^{(-j)}|}M\Big(\frac{\log n}{n}\Big)^{1/2}\;\geq\;\frac{M}{|\mathcal{N}_{i}^{(-j)}|}\Big(\frac{\log n}{n}\Big)^{1/2},$
which is strictly larger than $C_{2}(\log n/n)^{1/2}$ for $M>C_{2}$ and $|\mathcal{N}_{i}^{(-j)}|\geq 1$, a contradiction. Hence $|\mathcal{B}_{i}|=0$, and for every $k\in\mathcal{N}_{i}^{(-j)}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;M\Big(\frac{\log n}{n}\Big)^{1/2}.$
Applying Lemma 3.6 to each such $k$ yields that $U_{k}$ lies in the same block as $U_{i}$ and
$|U_{k}-U_{i}|\leq C^{\prime}r_{n}$
for some constant $C^{\prime}>0$. Since $U_{j}\in\mathcal{I}_{\varepsilon}$, the points $(U_{i},U_{j})$ and $(U_{k},U_{j})$ lie in the same Lipschitz region of $W$ in the first argument, so by Assumption 3.5(i),
$|P_{kj}-P_{ij}|=|W(U_{k},U_{j})-W(U_{i},U_{j})|\leq L|U_{k}-U_{i}|\leq LC^{\prime}r_{n}.$
Averaging over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$|B^{(-j)}_{ij}|\;\leq\;Cr_{n}$
for some $C>0$, as claimed. ∎
Lemma 3.7 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $B_{ij}^{(-j)}$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.

### v4

@@ 3.1 Neighborhood bias
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2}.$
Moreover, applying Theorem 1 of Zhang et al. (2017) to the (n1)(n-1)-node graph $A^{(-j)}$ yields, with probability at least $1-c_{1}n^{-c_{2}}$, uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Identifying ZLZ’s notation via
@@
the ZLZ bounds (with (n1)(n-1) in place of $n$) give
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
@@
The decomposition
$\widetilde{P}_{i\ell}-P_{i\ell}=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\bigl[(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\bigr]$
is identical to ZLZ equation (1), and the inequality
$(\widetilde{P}_{i\ell}-P_{i\ell})^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
follows exactly as in their argument, with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. Next, ZLZ’s control of two-hop quantities
$\frac{1}{n}(A^{2})_{ab}-\frac{1}{n}(P^{2})_{ab}$
is established in equations (4)-(6) of their supplementary material. Lemma 3.1 shows that the same bounds hold when $A$ is replaced by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, all subsequent ZLZ steps that rely on these concentration inputs remain valid for the leave-one-out graph.
ZLZ bound the row-average of $J_{1}$ using their decomposition (11) and bounds (12), (14), and (15). The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, yielding
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Similarly, ZLZ bound the row-average of $J_{2}$ using Cauchy-Schwarz and their “claim 2 of Lemma 2,” concluding in equation (16). Again, Lemma 3.1 ensures the same concentration inputs hold for $A^{(-j)}$, giving
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Combining these bounds yields the stated row-wise risk bound for the leave-one-out estimator. ∎
@@
Remark 3.6 (Scope of Assumption 3.5.).
This assumption excludes graphons with flat regions or non-identifiable row structure, and is tailored to settings where rowwise $L^{2}$ proximity reflects latent proximity within each block. In particular, it does not cover graphons with a growing number of blocks or highly irregular multiscale structure. The condition is standard in nonparametric graphon estimation when entrywise inference is desired, but we emphasize that it restricts the theory to graphons with fixed, well-separated smooth components.
Remark 3.7 (Good-node and good-edge sets.).
Fix $\varepsilon>0$ satisfying the feasibility condition
$b_{m}+\varepsilon<b_{m+1}-\varepsilon\qquad\text{for all }m=0,\dots,K-1,$
so that each interior interval $b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ is nonempty. Define the block interior
$I_{\varepsilon}:=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,\;b_{m+1}-\varepsilon].$
Let the random good-node set be
$\mathcal{G}_{\varepsilon}:=\{i\in[n]:\ \xi_{i}\in I_{\varepsilon}\},$
and the good-edge set
$\mathcal{E}_{\varepsilon}:=\{(i,j):i\neq j,\ i\in\mathcal{G}_{\varepsilon},\ j\in\mathcal{G}_{\varepsilon}\}.$
Since $\xi_{i}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}[0,1]$, we have $|\mathcal{G}_{\varepsilon}|\sim\mathrm{Bin}\big(n,\lambda(I_{\varepsilon})\big)$. In particular, for any $t>0$ Hoeffding’s inequality gives
$\mathbb{P}\!\left(\left|\frac{|\mathcal{G}_{\varepsilon}|}{n}-\lambda(I_{\varepsilon})\right|\geq t\right)\leq 2\exp(-2nt^{2}).$
Consequently, for fixed $\varepsilon>0$ the fraction of good nodes satisfies $|\mathcal{G}_{\varepsilon}|/n=\lambda(I_{\varepsilon})+O_{\mathbb{P}}(n^{-1/2})$, and if $\lambda(I_{\varepsilon})=1-O(\varepsilon)$ then $|\mathcal{G}_{\varepsilon}|=n(1-O(\varepsilon))$ with exponentially high probability.
@@
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Let $r_{n}=(\log n/n)^{1/4}$. There exist constants $c_{0},c_{1}>0$ depending only on Assumption 3.5 and $\varepsilon$ such that the following holds. If
$M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq c_{0},\qquad\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq M_{n}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},$
then with probability at least $1-n^{-c}$ (for some $c>0$) and for all sufficiently large $n$,
$U_{i},U_{k}\text{ lie in the same block},\qquad|U_{i}-U_{k}|\leq c_{1}\sqrt{M_{n}}\,r_{n}.$
@@
Let $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$, and define
$D(U_{i},U_{k})=\int_{0}^{1}(f_{i}(x)-f_{k}(x))^{2}\,dx.$
Concentration of the empirical row distance. Conditional on $(U_{i},U_{k})$, the variables
$Z_{\ell}:=(P_{i\ell}-P_{k\ell})^{2}=(W(U_{i},U_{\ell})-W(U_{k},U_{\ell}))^{2}$
are i.i.d. in $\ell$, bounded in $[0,1]$, and satisfy $\mathbb{E}[Z_{\ell}\mid U_{i},U_{k}]=D(U_{i},U_{k})$. Hoeffding’s inequality gives, for any $t>0$,
$\mathbb{P}\!\left(\left|\frac{1}{n}\sum_{\ell=1}^{n}Z_{\ell}-D(U_{i},U_{k})\right|\geq t\ \middle|\ U_{i},U_{k}\right)\leq 2e^{-2nt^{2}}.$
Taking $t=C(\log n/n)^{1/2}$ yields that, with probability at least $1-n^{-c}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq D(U_{i},U_{k})-C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.5)
If $U_{i},U_{k}$ lie in different blocks, Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$. Using (3.5), for large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\geq\frac{\eta^{2}\delta}{2}.$
Choosing $c_{0}=\eta^{2}\delta/4$ makes this incompatible with the assumed upper bound. Thus $U_{i},U_{k}$ must lie in the same block.
On the event of $U_{i},U_{k}$ lie in the same block and belong to $\mathcal{I}_{\varepsilon}$, so Assumption 3.5(iii) implies
$D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}.$
Combining this with (3.5) and the assumed upper bound gives
$\kappa(\varepsilon)|U_{i}-U_{k}|^{2}\leq\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}+C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}\leq(M_{n}+C)\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Therefore,
$|U_{i}-U_{k}|\leq\sqrt{\frac{M_{n}+C}{\kappa(\varepsilon)}}\,\Bigl(\frac{\log n}{n}\Bigr)^{1/4}\leq c_{1}\sqrt{M_{n}}\,r_{n},$
after adjusting constants. ∎
Corollary 3.9 (Bias bound on the good set).
Fix $\varepsilon>0$ satisfying the feasibility condition above and suppose the neighborhood regularity condition (display (3.4) in the main text) holds with constant $C_{2}$. Then there exists a constant $C(\varepsilon,C_{2})>0$ (depending on $L,\eta,\delta,\kappa(\varepsilon)$ and $C_{2}$) such that, with probability at least $1-c_{1}n^{-c_{2}}$,
$\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|=O_{\mathbb{P}}\big(C(\varepsilon,C_{2})\,r_{n}\big),\qquad r_{n}=(\log n/n)^{1/4}.$
Assumption 3.10 (Neighborhood regularity).
For each $(i,j)\in\mathcal{E}_{\varepsilon}$, the leave-one-out neighborhood $\mathcal{N}_{i}^{(-j)}$ satisfies
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\lesssim\;\Bigl(\frac{\log n}{n}\Bigr)^{1/2}$
with probability at least $1-n^{-c}$.
Remark.
Assumption 3.10 is consistent with the rowwise risk bounds of Zhang-Levina-Zhu together with the two-hop stability lemmas established above. We do not provide a fully explicit derivation of this condition for LOO neighborhoods here, but it is compatible with existing theory.
@@
Assume Assumption 3.5 and let $r_{n}=(\log n/n)^{1/4}$. Suppose that, with probability at least $1-n^{-c}$, for all $i\neq j$,
$\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\leq C_{2}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.6)
Then for all $i\neq j$ with $U_{i}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\leq Cr_{n}.$
The canonical choice $h_{n}\asymp\sqrt{n\log n}$ balances the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$. Undersmoothing ($h_{n}=o(\sqrt{n\log n})$) reduces bias at the cost of larger variance, while oversmoothing increases bias. Thus the canonical scaling is optimal for first-order mean-squared error and yields first-order coverage up to the explicit bias cushion.
@@
Fix $i\neq j$ and write $f_{u}(x)=W(u,x)$. Let $N=N_{i}^{(-j)}$ and define
$D(u,u^{\prime})=\int_{0}^{1}(f_{u}-f_{u^{\prime}})^{2}.$
First we show that conditional on $(U_{i},U_{k})$, the variables $(P_{i\ell}-P_{k\ell})^{2}$ are i.i.d. in $\ell$, bounded in $[0,1]$, and have mean $D(U_{i},U_{k})$. Hoeffding’s inequality and a union bound over $k\in N$ (using $|N|\leq h_{n}\asymp\sqrt{n\log n}$) give that with probability $1-n^{-c}$,
$\sup_{k\in N}\left|\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}-D(U_{i},U_{k})\right|\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.7)
Combining (3.6) and (3.7),
$\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k})\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Now we control the average latent distance between $U_{i}$ and its neighbors in $N$. Let
$\mathcal{C}_{i}=\{k\in N:U_{k}\text{ lies in a different block than }U_{i}\}.$
Assumption 3.5(ii) gives $D(U_{i},U_{k})\geq\eta^{2}\delta$ for $k\in\mathcal{C}_{i}$. Using (3.8),
$\frac{|\mathcal{C}_{i}|}{|N|}\eta^{2}\delta\leq C^{\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2},\qquad\Rightarrow\qquad\frac{|\mathcal{C}_{i}|}{|N|}\leq C^{\prime\prime}\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$
Decompose
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}=\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}+\frac{1}{|N|}\sum_{k\in\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}.$
The second term is at most $|\mathcal{C}_{i}|/|N|$. For $k\in N\setminus\mathcal{C}_{i}$ (same block), Assumption 3.5(iii) gives $D(U_{i},U_{k})\geq\kappa(\varepsilon)|U_{i}-U_{k}|^{2}$, hence
$\frac{1}{|N|}\sum_{k\in N\setminus\mathcal{C}_{i}}|U_{i}-U_{k}|^{2}\leq\frac{1}{\kappa(\varepsilon)}\cdot\frac{1}{|N|}\sum_{k\in N}D(U_{i},U_{k}).$
Using (3.8) and $(\dagger)$,
$\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}\leq C\Bigl(\frac{\log n}{n}\Bigr)^{1/2}.$ (3.8)
Step 3: Bias bound. Lipschitz continuity in the first argument gives
$|P_{kj}-P_{ij}|\leq L|U_{k}-U_{i}|.$
Thus
$|B^{(-j)}_{ij}|=\left|\frac{1}{|N|}\sum_{k\in N}(P_{kj}-P_{ij})\right|\leq L\sqrt{\frac{1}{|N|}\sum_{k\in N}|U_{i}-U_{k}|^{2}}\leq Cr_{n},$
using (3.8) and $r_{n}=(\log n/n)^{1/4}$. ∎
Lemma 3.11 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $\sup_{\begin{subarray}{c}i\neq j:\\ (i,j)\in\mathcal{E}_{\varepsilon}\end{subarray}}|B_{ij}^{(-j)}|$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.
Remark 3.12.
Under the canonical choice $h_{n}\asymp\sqrt{n\log n}$, the stochastic fluctuation scale $h_{n}^{-1/2}$ and the bias scale $r_{n}=(\log n/n)^{1/4}$ are of the same order. Thus the normal approximation is first-order accurate, and bias cannot be ignored without undersmoothing. Choosing a smaller bandwidth would reduce bias at the cost of increased variance, while larger bandwidths inflate bias. The canonical scaling balances these two effects.